An IM injection of Tobramycin 2.5mg/kg/day Q8H is ordered for a child with septicemia. The medication is supplied as 40mg/1mL. The child weighs 44.88 lbs.

How many mL should the nurse administer per dose?
Answer #1

Ans) Order: An IM injection of Tobramycin 2.5mg/kg/day Q8H.

Child's weight = 2.5 × 44.88 = 112.2 mg/day

In three divided dose 8 hourly : 112.2/3 = 37.4 mg each dose.

Available: 40 mg/ 1ml

40 mg = 1 ml

37.4 mg = x

x = 37.4/40 = 0.934 ~ 0.9 ml

The nurse needs to administer 0.9 ml for each of the dose 8 hourly.